OpenStudy (anonymous):

Please tell me if I am correct? Six is the geometric mean of 18 and what number? 4 3 2 9 Is the right answer 2?

jimthompson5910 (jim_thompson5910):

If you have 2 numbers x and y, then the geometric mean of those two numbers is G = sqrt(x*y)

jimthompson5910 (jim_thompson5910):

In this case, we're given x = 18 y = unknown (leave it as y for now) G = 6

jimthompson5910 (jim_thompson5910):

so G = sqrt(x*y) 6 = sqrt(18y) 6^2 = (sqrt(18y))^2 36 = 18y 18y =36 y = ???

OpenStudy (anonymous):

So it would be 2 ?

jimthompson5910 (jim_thompson5910):

correct